OpenStudy (anonymous):

Consider the graph of the cosine function shown below. a. Find the period and amplitude of the cosine function. b. At what values of for do the maximum value(s), minimum values(s), and zeros occur?

OpenStudy (jdoe0001):

the cosine function "shown below" looks blank, so I guess it has no period or amplitude or anything :/

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Amplitude is 4 units & Period is Pi

OpenStudy (anonymous):

Maxima Occurs at n*pi where n =1,2,3,4....... & so-on Minima Occurs at n*pi/2 where n =1,2,3,4....... & so-on Zero Occurs at n*pi/4 where n =1,2,3,4....... & so-on

OpenStudy (anonymous):

thank you :)